OpenStudy (anonymous):

1 3/4 + 8 1/3 + 7 4/5 ? must be interger or mixed number. simplify please

OpenStudy (anonymous):

(7/4)+(25/3)+(39/5)

OpenStudy (anonymous):

find common denominator

OpenStudy (anonymous):

i cant come up with one

OpenStudy (anonymous):

3*4*5=?

OpenStudy (anonymous):

60

OpenStudy (anonymous):

theres ur common denominator

OpenStudy (anonymous):

so then i do 1 3/60 + 8 1/60 + 7 4/60?

OpenStudy (anonymous):

no u need to find a number that if u multiplied it by 4 u get 60 so 4x__=60 what goes in the blank?

OpenStudy (amistre64):

we was taught to line them up in my day 1 3/4 8 1/3 7 4/5 ------ 16+N \[\frac{3}{4}+\frac{1}{3}+\frac{4}{5} = N \text{ ; *20}\] \[\frac{3(20)}{4}+\frac{1(20)}{3}+\frac{4(20)}{5} = 20N \text{ ; reduce}\] \[\frac{3(5)}{1}+\frac{1(20)}{3}+\frac{4(4)}{1} = 20N \text{ ; *3}\] \[\frac{3(5)(3)}{1}+\frac{1(20)(3)}{3}+\frac{4(4)(3)}{1} = 20(3)N \text{ ; reduce}\] \[\frac{3(5)(3)}{1}+\frac{1(20)(1)}{1}+\frac{4(4)(3)}{1} = 60N \text{ ; re-write}\] \[45+20+36 = 60N \text{ ; add}\] \[101 = 60N \text{ ; divide of the 60}\] \[\frac{101}{60} = N \text{ ; divide of the 60}\]

OpenStudy (amistre64):

change N to a mixed numeral by dividing it out

OpenStudy (anonymous):

so the answer is 1 41/60?

OpenStudy (amistre64):

16 + N 16 + (1 + 41/60) = 17' 41/60 if you did the division right ;)

OpenStudy (anonymous):

so he answer is 17 41/60?

OpenStudy (anonymous):

the

OpenStudy (amistre64):

id say it is

OpenStudy (anonymous):

it says that it is incorrect

OpenStudy (amistre64):

then we will need to see why ...

OpenStudy (anonymous):

how would be do that?

OpenStudy (amistre64):

17.88333333 S = .883333... 1000S = 883.3333 -100S = 88.33333 ----------------- 900S = 795 S = 795/900 = 53/60

OpenStudy (amistre64):

12*4 = 48; not 36 .. i forgot how to multiply

OpenStudy (akshay_budhkar):

lol amistre

OpenStudy (amistre64):

:) so N = 113/60

OpenStudy (akshay_budhkar):

Legend!!!